OpenStudy (anonymous):

Please Help! I will fan and medal. Use the inverse properties of logarithms to simplify the expression. e^(ln 3)

jimthompson5910 (jim_thompson5910):

Rule: \[\LARGE e^{\ln(x)} = x\] \[\LARGE \ln\left(e^x\right) = x\]

OpenStudy (anonymous):

Is that the simplified equation? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you will use one of those two equations to answer the question

OpenStudy (anonymous):

how do i know which one

jimthompson5910 (jim_thompson5910):

the question is e^(ln(3)) there's only one number in it: 3 so why not replace x with 3

jimthompson5910 (jim_thompson5910):

and then try to match up the question with one of the equations given above

OpenStudy (anonymous):

Okay so it would be the first equation? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

So what now

jimthompson5910 (jim_thompson5910):

replace x with 3

OpenStudy (anonymous):

e^(ln(3))=3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

the e^x and ln(x) functions are inverses of each other one goes forward, the other takes you in reverse so they undo each other it's like multiplication and division

OpenStudy (anonymous):

alright so now what? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're done. The answer is 3

jimthompson5910 (jim_thompson5910):

\[\LARGE e^{\ln(x)} = x\] \[\LARGE e^{\ln(3)} = 3\]

jimthompson5910 (jim_thompson5910):

\(\LARGE e^{\ln(3)}\) simplifies to \(\LARGE 3\)

OpenStudy (anonymous):

so the answer is 3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome